OpenStudy (anonymous):

Help for MEDAL : question about sequences n 1 2 3 4 5 6 f(n) 1 2 3 5 8 13

OpenStudy (anonymous):

Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

@briannamooney

OpenStudy (anonymous):

can you help @shinalcantara

OpenStudy (anonymous):

you were right why did you delete your answer

OpenStudy (shinalcantara):

i found it having problem when having it as a function of 5 & 6.. there's something wrong about it.. it will only be true to 1,2,3 & 4

OpenStudy (anonymous):

ohh but it was still the 2nd one